OpenStudy (anonymous):

The sum of Jack and his father's age is 52. Jack's father's age is 2 less than 5 times Jack's age. Find the ages of Jack and his father.

OpenStudy (anonymous):

Jack is 9 Jack's father is 43

OpenStudy (anonymous):

(5x-2)+x=52 6x =52 +2 x= 54/6 x=9 (9*5)-2= 43 Jack is 9 his father is 43